# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ B.5.2 Polynomial interpolation on infinite grid
$\|f_{i}^{\prime}\|_{1}\leq(4c_{L}+1)^{d}\|f^{\prime}\|_{\infty}\lesssim\|f\|_{\infty}.$ (74)
Furthermore, in view of (73) and (74) and noting that in view of the finite co-domain $\ell^{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})=\ell^{\infty}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})$,
$\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}\|_{\infty}=$ $\sup_{f\in\ell^{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})}\frac{\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)\|_{\infty}}{\|f\|_{\infty}}$
$=$ $\sup_{f\in\ell^{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})}\frac{\sup_{X_{i}\in\tilde{\mathcal{X}}_{n}}|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(X_{i})|}{\|f\|_{\infty}}$
$\lesssim$ $\sup_{f_{i}^{\prime}\in\ell^{1}(\mathbb{Z}^{d})}\sup_{X_{i}\in\tilde{\mathcal{X}}_{n}}\frac{|P^{j_{2}-j_{3}}(f_{i}^{\prime})(2^{j_{2}+1}X_{i})|}{\|f_{i}^{\prime}\|_{1}}$
$\leq$ $\sup_{j\in\mathbb{N}}\sup_{f_{i}^{\prime}\in\ell^{1}(\mathbb{Z}^{d})}\frac{\|P^{j}(f_{i}^{\prime})\|_{\infty}}{\|f^{\prime}_{i}\|_{1}}.$ (75)
Assertion (B.5.2) basically bounds the $\ell^{\infty}$ norm of the recursive polynomial interpolation operator on a finite grid with the supremum norm of the recursive polynomial interpolation operator on the infinite grid. From now on we will focus on bounding the latter term.
B.5.3 Bounding the interpolation operator via Fourier transform
The polynomial interpolation operator $P$ on the infinite grid is also a convolution operator, which motivates us to use Fourier analysis. For all $f\in\ell^{1}(\mathbb{Z}^{d})$, its (discrete time) Fourier transform is denoted by $\mathcal{F}:\ell^{1}(\mathbb{Z}^{d})\to\ell^{1}(\mathbb{T}([0,1]^{d}))$, with $\mathbb{T}([0,1]^{d})$ denoting the $d$-dimensional torus, such that for all $\xi\in\mathbb{T}([0,1]^{d})$,
$\mathcal{F}(f)(\xi)=\sum_{z\in\mathbb{Z}^{d}}e^{-2\pi\iota\langle z,\xi\rangle}f(z).$ (76)
Furthermore, for $\rho\in\ell^{1}(\mathbb{T}([0,1]^{d}))$ and $z\in\mathbb{Z}^{d}$, we also define the inverse (discrete time) Fourier transform as
$\mathcal{F}^{-1}(\rho)(z)=\int_{\xi\in\mathbb{T}([0,1]^{d})}e^{2\pi\iota\langle z,\xi\rangle}\rho(\xi)d\xi.$ (77)
For all $f\in\ell^{1}(\mathbb{Z}^{d})$, we have the following identity
@@ B.5.3 Bounding the interpolation operator via Fourier transform
We proceed with proving Lemma 14.
Proof of Lemma 14.
In view of the analysis above, it suffices to bound the right hand side term in (B.5.2). The proof is divided into two parts. First we consider the special case where $d=1$ and $m=\underline{\alpha}+1$ is an even integer. We demonstrate in this simpler case the main ideas of the proof. We show that the polynomial interpolation operator $P$ acting on the function $f$ in the Fourier domain can be expressed as a Markov operator acting on $\mathcal{F}(f)$. Building on this assertion we derive the stated upper bound for the recursive application of $P$. Then we consider the rest of the cases. Here, in the more general setting, the previous technique can not be applied directly, but one has to introduce additional technical steps.
Case 1: We first consider the case $d=1$ and $m=\underline{\alpha}+1$ an even integer. Then the multi-index set $\mathbb{K}$, see (68), has only a single element $k=1$ and the parent set of $k/2$ is $A=\{-m/2+1,-m/2+2,\cdots,m/2-1,m/2\}$. The operator $L_{1}$, given in (69) for $k=1$, is
$L_{1}(f)(z)=v_{1/2}^{T}V_{A}^{-1}\bm{f}(z+A).$
@@
$\mathcal{F}(L_{1}(f))(\xi)=\sum_{j=1}^{m/2}b_{j}(e^{2j\pi\iota\xi}+e^{2(-j+1)\pi\iota\xi})\mathcal{F}(f)(\xi),\;\;\forall\xi\in\mathbb{T}([0,1]).$
Then, in view of the definition of the operator $P$ in (70), the Fourier transform of $P(f)$ can be computed as
@@
$=$ $\mathcal{F}(f)(2\xi)+e^{-2\pi\iota\xi}\mathcal{F}(L_{1}(f))(2\xi)$
@@
Let us denote the above linear operator by $M:\ell^{1}(\mathbb{T}([0,1]))\to\ell^{1}(\mathbb{T}([0,1]))$, i.e.
$\mathcal{F}(P(f))=M(\mathcal{F}(f)).$ (78)
@@
$\|M\rho\|_{1}=\|\rho\|_{1},\forall\rho\in\ell^{1}(\mathbb{T}).$ (79)
By definition, $\mathcal{F}(P(f))$ is the Fourier transform of the polynomial interpolation of the function $f$. Thus, $\sum_{j=1}^{m/2}b_{j}2\cos(2\pi(2j-1)\xi)$ is the frequency response of polynomial interpolation in the literature of signal processing, which has complex modulus no greater than one (page 42 of [31]). As a result, $M\rho$ is nonnegative as long as $\rho$ is. It remains to show that $M\rho$ integrates to one. We directly compute the integral as
@@
Then, for all $f$, define $\mathcal{F}(f)_{+}(\xi)=\max\{\mathcal{F}(f)(\xi),0\}$ and $\mathcal{F}(f)_{-}(\xi)=\mathcal{F}(f)(\xi)-\mathcal{F}(f)_{+}(\xi)$. Furthermore, define $f_{+}=\mathcal{F}^{-1}(\mathcal{F}(f)_{+})$ and $f_{-}=f-f_{+}$. Combining equations (78) and (79), we have
@@
where in the last line, $\|\mathcal{F}(f)\|_{1}\leq\|\mathcal{F}(f)\|_{\infty}$ follows from the fact that $\mathcal{F}(f)$ is defined on torus $\mathbb{T}([0,1]^{d})$. The inequalities $\|\mathcal{F}(f)\|_{\infty}\leq\|f\|_{1}$ and $\|P^{j}(f)\|_{\infty}\leq\|\mathcal{F}(P^{j}(f))\|_{1}$ follow from the definition of Fourier transform (76) and (77), respectively. Therefore, the right hand side of (B.5.2) is bounded from above by one, which in turn implies that $\vartheta_{n}=O(1)$.
Case 2: Next we extend the results to the more general case of $d$ and $\alpha$. The proof is further divided into two steps. The first step is similar to Case 1. We study the operator $P$ in Fourier domain, show that it is a linear operator $M$ and find the associated kernel $\mathcal{K}_{M}$. However, the operator $M$ is not necessarily Markov in this general setting and hence can not be directly used to derive the stated upper bounds. Therefore, we introduce a second step, by converting the operator $M$ into another operator $\tilde{M}$, which will be used in the analysis.
Step 1 Recall the multi-index set $\mathbb{K}$ defined in (68), and for $k\in\mathbb{K}$, denote the vector
@@
Denote the parent set of $k/2$ as $\mathrm{pa}(k/2)=\{z_{k,1},z_{k,2},\cdots,z_{k,m}\}$, which is a corner set as in Section 4.2 with $m={\underline{\alpha}+d\choose\underline{\alpha}}$. Then the discrete time Fourier transform of $L_{k}(f)$, given in (69), is
$\mathcal{F}(L_{k}(f))(\xi)=\sum_{k=1}^{m}b_{k,m}e^{2\pi\iota\langle z_{k,m},\xi\rangle}\mathcal{F}(f)(\xi).$
In view of the definition of the operator $P$ in (70),
@@
$=$ $\sum_{z\in\mathbb{Z}^{d}}f(z)e^{-2\pi\iota\langle z,2\xi\rangle}+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}\sum_{z\in\mathbb{Z}^{d}}L_{k}(f)(z)e^{-2\pi\iota\langle z+k/2,2\xi\rangle}$
$=$ $\mathcal{F}(f)(2\xi)+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}e^{-2\pi\iota\langle k,\xi\rangle}\mathcal{F}(L_{k}(f))(2\xi)$
$=$ $\mathcal{F}(f)(2\xi)\left(1+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}e^{-2\pi\iota\langle k,\xi\rangle}\sum_{k=1}^{m}b_{k,m}e^{4\pi\iota\langle z_{k,m},\xi\rangle}\right).$ (80)
@@
and a linear operator $M:\ell^{1}(\mathbb{T}([0,1]^{d}))\to\ell^{1}(\mathbb{T}([0,1]^{d})$ as
@@
Then the equation (B.5.3) can be equivalently written as
$\mathcal{F}(P(f))=M(\mathcal{F}(f)).$ (81)
@@
$\mathcal{K}_{M}(\xi_{1},A)=\frac{1}{2^{d}}\sum_{\xi_{2}\in\mathbb{T}([0,1]^{d}):\,2\xi_{2}=\xi_{1}}\psi(\xi_{2})\mathbbm{1}_{\xi_{2}\in A}.$ (82)
Next we show that the above kernel $\mathcal{K}_{M}$ is associated to the linear operator $M$. It is easy to see that for all $\xi_{1}\in\mathbb{T}([0,1]^{d})$, $\mathcal{K}_{M}(\xi_{1},\cdot)$ defines an $\ell^{1}$ function on $\mathbb{T}([0,1]^{d})$ that integrates to one. Moreover, letting the multi-index set $\mathbb{K}_{0}$ be
@@
$=$ $\int_{A}\rho(2\xi_{2})\psi(\xi_{2})d\xi_{2}=\int_{A}M(\rho)(\xi_{2})d\xi_{2}.$
hence $\mathcal{K}_{M}$ is indeed the kernel associated with the operator $M$. In particular, for $A=\mathbb{T}([0,1]^{d})$,
$\int_{\mathbb{T}([0,1]^{d})}M(\rho)(\xi)d\xi=\int_{\mathbb{T}([0,1]^{d})}\rho(\xi)\mathcal{K}_{M}(\xi,\mathbb{T}([0,1]^{d}))d\xi.\\$ (83)
Step 2 Note, that in view of assertion (83), if $\mathcal{K}_{M}(\xi,\cdot)$ would be a probability measure, then the same lines of reasoning apply as in Case 1. However, $\mathcal{K}_{M}(\xi,\cdot)$ may output complex values by its definition. To overcome this problem, we transform $M$ into another operator whose associated kernel provides finite, non-negative measures.
First, for $\xi\in\mathbb{T}([0,1]^{d})$, define the vector $\ell^{1}$ norm on the torus as
$\|\xi\|_{\ell^{1}}=\sum_{j=1}^{d}\min\{\xi[j],1-\xi[j]\}.$
Note that the $\ell^{1}$ norm of $\xi$ is the $\ell^{1}$ distance between $\xi$ and the origin on torus. Furthermore, we define the operator $\Xi:\ell^{1}(\mathbb{T}([0,1]^{d}))\to\ell^{1}(\mathbb{T}([0,1]^{d}))$ such that for all $\xi\in\mathbb{T}([0,1]^{d})$,
$\Xi(\rho)(\xi)=\|\xi\|_{\ell^{1}}|\rho(\xi)|.$
Then, for any function $f\in\ell^{1}(\mathbb{Z}^{d})$ and for all $z\in\mathbb{Z}^{d}$,
@@
By the definition of inverse Fourier transform (77), $\mathcal{F}(f)$ is an $\ell^{\infty}$ function on $\mathbb{T}([0,1]^{d})$. Therefore, we can extend the domain of the function $f$ from $\mathbb{Z}^{d}$ to $\mathbb{R}^{d}$ as
@@
such that $f_{\mathrm{ext}}(z)=f(z),\forall z\in\mathbb{Z}^{d}$ and $f_{\mathrm{ext}}\in\ell^{\infty}(\mathbb{R}^{d})$. Moreover, $f_{\mathrm{ext}}$ has uniformly bounded gradient, i.e. for all $x\in\mathbb{R}^{d}$
@@
$\leq$ $2\pi\sum_{k\in\mathbb{N}^{d}:0\leq k[j]\leq 1}\int_{[0,1/2]^{d}+k/2}\|\xi\|_{\ell^{1}}|\mathcal{F}(f)(\xi)|d\xi$
$=$ $2\pi\int_{\mathbb{T}([0,1]^{d})}\|\xi\|_{\ell^{1}}|\mathcal{F}(f)(\xi)|d\xi$

### v4

@@ E.5.2 Polynomial interpolation on infinite grid
$\|f_{i}^{\prime}\|_{1}\leq(4c_{L}+1)^{d}\|f^{\prime}\|_{\infty}\lesssim\|f\|_{\infty}.$ (E.28)
Furthermore, by recalling (E.27), (E.28) and noting that in view of the finite co-domain $L_{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})=L_{\infty}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})$, we have
$\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}\|=$ $\sup_{f\in L_{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})}\frac{\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)\|_{\infty}}{\|f\|_{\infty}}$
$=$ $\sup_{f\in L_{1}(\cup_{j=0}^{j_{1}}\tilde{\mathcal{N}}_{j})}\frac{\sup_{X_{i}\in\tilde{\mathcal{X}}_{n}}|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(X_{i})|}{\|f\|_{\infty}}$
$\lesssim$ $\sup_{f_{i}^{\prime}\in L_{1}(\mathbb{Z}^{d})}\sup_{X_{i}\in\tilde{\mathcal{X}}_{n}}\frac{|P^{j_{2}-j_{3}}(f_{i}^{\prime})(2^{j_{2}+1}X_{i})|}{\|f_{i}^{\prime}\|_{1}}$
$\leq$ $\sup_{j\in\mathbb{N}}\sup_{f_{i}^{\prime}\in L_{1}(\mathbb{Z}^{d})}\frac{\|P^{j}(f_{i}^{\prime})\|_{\infty}}{\|f^{\prime}_{i}\|_{1}}.$ (E.29)
Assertion (E.5.2) basically bounds the $L_{\infty}$ norm of the recursive polynomial interpolation operator on a finite grid with the supremum norm of the recursive polynomial interpolation operator on the infinite grid. From now on we will focus on bounding the latter term.
E.5.3 Bounding the interpolation operator via Fourier transform
The polynomial interpolation operator $P$ on the infinite grid is also a convolution operator, which motivates us to use Fourier analysis. For all $f\in L_{1}(\mathbb{Z}^{d})$, its (discrete time) Fourier transform is denoted by $\mathcal{F}:L_{1}(\mathbb{Z}^{d})\to L_{1}(\mathbb{T}([0,1]^{d}))$, with $\mathbb{T}([0,1]^{d})$ denoting the $d$-dimensional torus, such that for all $\xi\in\mathbb{T}([0,1]^{d})$,
$\mathcal{F}(f)(\xi)=\sum_{z\in\mathbb{Z}^{d}}e^{-2\pi\iota\langle z,\xi\rangle}f(z).$ (E.30)
Furthermore, for $\rho\in L_{1}(\mathbb{T}([0,1]^{d}))$ and $z\in\mathbb{Z}^{d}$, we also define the inverse (discrete time) Fourier transform as
$\mathcal{F}^{-1}(\rho)(z)=\int_{\xi\in\mathbb{T}([0,1]^{d})}e^{2\pi\iota\langle z,\xi\rangle}\rho(\xi)d\xi.$ (E.31)
For all $f\in L_{1}(\mathbb{Z}^{d})$, we have the following identity
@@ E.5.3 Bounding the interpolation operator via Fourier transform
We now proceed to the proof of Lemma E.1.
Proof of Lemma E.1.
In view of the analysis above, it suffices to bound the right-hand side term in (E.5.2). The proof is divided into two parts. First, we consider the special case where $d=1$ and $m=\underline{\alpha}+1$ is an even integer. This simpler case illustrates the main ideas of the proof. We show that the polynomial interpolation operator $P$, acting on the function $f$ in the Fourier domain, can be expressed as a Markov operator acting on $\mathcal{F}(f)$. Building on this observation, we derive the stated upper bound for the recursive application of $P$. In the second part, we address the more general cases. Here, the previous technique cannot be applied directly, and additional technical steps are required.
Case 1: We first consider the case $d=1$ and $m=\underline{\alpha}+1$ an even integer. Then the multi-index set $\mathbb{K}$, see (E.22), has only a single element $k=1$ and the parent set of $k/2$ is $A=\{-m/2+1,-m/2+2,\cdots,m/2-1,m/2\}$. The operator $\mathrm{In}_{1}$, given in (E.23) for $k=1$, is
$\mathrm{In}_{1}(f)(z)=v_{1/2}^{T}V_{A}^{-1}\bm{f}(z+A).$
@@
$\mathcal{F}(\mathrm{In}_{1}(f))(\xi)=\sum_{j=1}^{m/2}b_{j}(e^{2j\pi\iota\xi}+e^{2(-j+1)\pi\iota\xi})\mathcal{F}(f)(\xi),\;\;\forall\xi\in\mathbb{T}([0,1]).$
Then, in view of the definition of the operator $P$ in (E.24), the Fourier transform of $P(f)$ can be computed as
@@
$=$ $\mathcal{F}(f)(2\xi)+e^{-2\pi\iota\xi}\mathcal{F}(\mathrm{In}_{1}(f))(2\xi)$
@@
Let us denote the above linear operator by $M:L_{1}(\mathbb{T}([0,1]))\to L_{1}(\mathbb{T}([0,1]))$, i.e.
$\mathcal{F}(P(f))=M(\mathcal{F}(f)).$ (E.32)
@@
$\|M\rho\|_{1}=\|\rho\|_{1},\forall\rho\in L_{1}(\mathbb{T}).$ (E.33)
By definition, $\mathcal{F}(P(f))$ is the Fourier transform of the polynomial interpolation of the function $f$. Thus, $\sum_{j=1}^{m/2}b_{j}2\cos(2\pi(2j-1)\xi)$ corresponds to the frequency response of polynomial interpolation in the literature of signal processing, whose complex modulus does not exceed one (see page 42 of [32]). Thus $M\rho\geq 0$ whenever $\rho\geq 0$. It remains to show that $M\rho$ integrates to one. We directly compute the integral as
@@
Then, for all $f$, define $\mathcal{F}(f)_{+}(\xi)=\max\{\mathcal{F}(f)(\xi),0\}$ and $\mathcal{F}(f)_{-}(\xi)=\mathcal{F}(f)(\xi)-\mathcal{F}(f)_{+}(\xi)$. Furthermore, define $f_{+}=\mathcal{F}^{-1}(\mathcal{F}(f)_{+})$ and $f_{-}=f-f_{+}$. Combining equations (E.32) and (E.33), we have
@@
where in the last line, $\|\mathcal{F}(f)\|_{1}\leq\|\mathcal{F}(f)\|_{\infty}$ follows from the fact that $\mathcal{F}(f)$ is defined on the torus $\mathbb{T}([0,1]^{d})$. The inequalities $\|\mathcal{F}(f)\|_{\infty}\leq\|f\|_{1}$ and $\|P^{j}(f)\|_{\infty}\leq\|\mathcal{F}(P^{j}(f))\|_{1}$ follow from the definition of Fourier transform (E.30) and (E.31), respectively. Therefore, the right-hand side of (E.5.2) is bounded from above by one, which in turn implies that $\vartheta_{n}=O(1)$.
Case 2: Next we extend the results to the more general case of $d$ and $\alpha$. The proof is divided into two steps. In the first step, analogous to Case 1, we study the operator $P$ in the Fourier domain, showing that it corresponds to a linear operator $M$ and identifying the associated kernel $\mathcal{K}_{M}$. However, in this general setting, $M$ is not necessarily a Markov operator and therefore cannot be directly used to derive the stated upper bounds. In the second step we address this issue by constructing a modified operator $\tilde{M}$, which is then employed in the analysis.
Step 1 Recall the multi-index set $\mathbb{K}$ defined in (E.22), and for $k\in\mathbb{K}$, introduce the vector
@@
Denote the parent set of $k/2$ as $\mathrm{pa}(k/2)=\{z_{k,1},z_{k,2},\cdots,z_{k,m}\}$, which is a corner set as in Section 3.3 with $m={\underline{\alpha}+d\choose\underline{\alpha}}$. Then the discrete time Fourier transform of $\mathrm{In}_{k}(f)$, given in (E.23), is
$\mathcal{F}(\mathrm{In}_{k}(f))(\xi)=\sum_{k=1}^{m}b_{k,m}e^{2\pi\iota\langle z_{k,m},\xi\rangle}\mathcal{F}(f)(\xi).$
In view of the definition of the operator $P$ in (E.24),
@@
$=$ $\sum_{z\in\mathbb{Z}^{d}}f(z)e^{-2\pi\iota\langle z,2\xi\rangle}+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}\sum_{z\in\mathbb{Z}^{d}}\mathrm{In}_{k}(f)(z)e^{-2\pi\iota\langle z+k/2,2\xi\rangle}$
$=$ $\mathcal{F}(f)(2\xi)+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}e^{-2\pi\iota\langle k,\xi\rangle}\mathcal{F}(\mathrm{In}_{k}(f))(2\xi)$
$=$ $\mathcal{F}(f)(2\xi)\left(1+\sum_{k:|k|\geq 1,0\leq k[j]\leq 1}e^{-2\pi\iota\langle k,\xi\rangle}\sum_{k=1}^{m}b_{k,m}e^{4\pi\iota\langle z_{k,m},\xi\rangle}\right).$ (E.34)
@@
and a linear operator $M:L_{1}(\mathbb{T}([0,1]^{d}))\to L_{1}(\mathbb{T}([0,1]^{d})$ as
@@
Then equation (E.5.3) can be reformulated as
$\mathcal{F}(P(f))=M(\mathcal{F}(f)).$ (E.35)
@@
$\mathcal{K}_{M}(\xi_{1},A)=\frac{1}{2^{d}}\sum_{\xi_{2}\in\mathbb{T}([0,1]^{d}):\,2\xi_{2}=\xi_{1}}\psi(\xi_{2})\mathbbm{1}_{\xi_{2}\in A}.$ (E.36)
Next, we show that the above kernel $\mathcal{K}_{M}$ is associated to the linear operator $M$. It is easy to see that for all $\xi_{1}\in\mathbb{T}([0,1]^{d})$, $\mathcal{K}_{M}(\xi_{1},\cdot)$ defines an $L_{1}$ function on $\mathbb{T}([0,1]^{d})$ that integrates to one. Moreover, letting the multi-index set $\mathbb{K}_{0}$ be
@@
$=$ $\int_{A}\rho(2\xi_{2})\psi(\xi_{2})d\xi_{2}=\int_{A}M(\rho)(\xi_{2})d\xi_{2},$
hence, $\mathcal{K}_{M}$ is indeed the kernel associated with the operator $M$. In particular, for $A=\mathbb{T}([0,1]^{d})$,
$\int_{\mathbb{T}([0,1]^{d})}M(\rho)(\xi)d\xi=\int_{\mathbb{T}([0,1]^{d})}\rho(\xi)\mathcal{K}_{M}(\xi,\mathbb{T}([0,1]^{d}))d\xi.\\$ (E.37)
Step 2 Note that, in view of assertion (E.37), if $\mathcal{K}_{M}(\xi,\cdot)$ were a probability measure, then the same reasoning as in Case 1 would apply. However, $\mathcal{K}_{M}(\xi,\cdot)$ may take complex values. To address this issue, we transform $M$ into a new operator whose associated kernel provides finite, non-negative measures.
First, for $\xi\in\mathbb{T}([0,1]^{d})$, define the vector $L_{1}$ norm on the torus as
$\|\xi\|_{L_{1}}=\sum_{j=1}^{d}\min\{\xi[j],1-\xi[j]\}.$
Note that the $L_{1}$ norm of $\xi$ is the $L_{1}$ distance between $\xi$ and the origin on torus. Furthermore, we define the operator $\Xi:L_{1}(\mathbb{T}([0,1]^{d}))\to L_{1}(\mathbb{T}([0,1]^{d}))$ such that for all $\xi\in\mathbb{T}([0,1]^{d})$,
$\Xi(\rho)(\xi)=\|\xi\|_{L_{1}}|\rho(\xi)|.$
Then, for any function $f\in L_{1}(\mathbb{Z}^{d})$ and for all $z\in\mathbb{Z}^{d}$,
@@
By the definition of the inverse Fourier transform (E.31), $\mathcal{F}(f)$ is an $L_{\infty}$ function on $\mathbb{T}([0,1]^{d})$. Therefore, we can extend the domain of the function $f$ from $\mathbb{Z}^{d}$ to $\mathbb{R}^{d}$ as
@@
such that $f_{\mathrm{ext}}(z)=f(z),\forall z\in\mathbb{Z}^{d}$ and $f_{\mathrm{ext}}\in L_{\infty}(\mathbb{R}^{d})$. Moreover, $f_{\mathrm{ext}}$ has uniformly bounded gradient, i.e. for all $x\in\mathbb{R}^{d}$
@@
$\leq$ $2\pi\sum_{k\in\mathbb{N}^{d}:0\leq k[j]\leq 1}\int_{[0,1/2]^{d}+k/2}\|\xi\|_{L_{1}}|\mathcal{F}(f)(\xi)|d\xi$
$=$ $2\pi\int_{\mathbb{T}([0,1]^{d})}\|\xi\|_{L_{1}}|\mathcal{F}(f)(\xi)|d\xi$